OpenStudy (anonymous):

3(sqrt -9)*3(sqrt -81)

OpenStudy (anonymous):

the answer is 81n

OpenStudy (anonymous):

ur welcome :))))

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

4 (\sqrt 125m)*4 (\sqrt 35)\]

OpenStudy (anonymous):

do you know how to get that tho?? cuz if you dont you could get into trouble on tests and stuff

OpenStudy (anonymous):

all you do is find the square roots first.......

OpenStudy (anonymous):

11.18m and 5.91

OpenStudy (anonymous):

11.18 *4 = 44.72m

OpenStudy (anonymous):

divide by four you just get 11.18 again divide that by5.91.......

OpenStudy (anonymous):

wich is 1.90m

OpenStudy (anonymous):

which is your answer.

OpenStudy (anonymous):

that wasnt any of choices but thanks for trying to help me.

OpenStudy (anonymous):

what were the choices?

OpenStudy (anonymous):

25 sqrt 7m 5m 4 sqrt 7 5 4sqrt 35m 5 4sqrt 7m

OpenStudy (anonymous):

what do u think sqrt stands for?? (i think im mixed up) it stands 4 square root right?